A skier starts from rest and slides 9.0m down a slope in 3.0s. In what time after starting will the skier acquire a speed of 24m/s? Assume that the acceleration is constant.


1
Expert's answer
2021-03-29T09:01:41-0400

The distance "h = 9m" covered by the skier under a constant acceleration from the rest is given by the kinematic equation:


"h =\dfrac{at_0^2}{2}"

where "a" is the acceleration and "t_0 = 3s" is the time of covering "h". Thus, obtain:


"a = \dfrac{2h}{t_0^2}"

On the other hand, by definition, the acceleration is:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = 24m\/s,v_i = 0m\/s" are the final and initial speeds respectively, and "t" is the time when the skier acquire a speed "v_f". Expressing "t" and substituting "a" obtain:


"t = \dfrac{v_f}{a} = \dfrac{v_ft_0^2}{2h}\\
t = \dfrac{24\cdot 3^2}{2\cdot 9} = 12s"

Answer. 12s.